# SA-ResGS: Self-Augmented Residual 3D Gaussian Splatting for Next Best View Selection

### v1

@@ 4 Experimental Results
We evaluate our approach on two benchmark datasets: NeRF-Synthetic nerf, and Mip-NeRF 360 (mip-nerf360). While both datasets comprise scenes ranging from synthetic object-scale to real-world outdoor environments with full 360-degree coverage, its uniform and curated camera trajectories provide limited challenge for active view selection, since even simple heuristics (e.g., furthest-distance selection) already perform reliably under balanced coverage (Xiao:CVPR24:NeRFDirector). To address this limitation, we carefully curate an extended benchmark dataset including seven diverse scenes from Deep Blending (deepblending) and Tanks and Temples (tankandtemples), which introduce unbalanced view distributions and varied scene scales that better reflect practical conditions. All experiments are conducted using images at their original resolutions, and further details on dataset curation are provided in Sec. A of the Appendix.
@@
We compare our method quantitatively and qualitatively against several active 3DGS baselines which operate solely on RGB images: FisherRF (fisherrf), ACP (acp), and random view selection. We also include 2D-based view selection methods adopted from Active View Selector framework (wang2024avs). Following this framework, we incorporated two image quality assessment (IQA) models (MUSIQ and CrossScore) to evaluate perceptual quality. Both models were re-implemented according to the authors’ official instructions and publicly available code.
@@ 4.1 Active View Selection
Table 1: Quantitative results for the Active View Selection. We compare our model with (1) Rule-based models (Random, ACP (acp)), (2) 2D-based models (MUSIQ, CrossScore (chen2024gennbv)), and 3D-based models (FisherRF (fisherrf)). Results are averaged over 9 scenes from the Mip-NeRF 360, 7 scenes from NeRF-synthetic dataset and 7 additional scenes from the Deep Blending and Tanks and Temples. We conduct four trials for each scene and report average scores. For statistics for all trial please refer to the Appendix E.
@@
Following the experimental protocols outlined by fisherrf, we adopt their prescribed initial view configurations and view selection schedules. Specifically, our experiments initiate with four uniformly distributed views, subsequently selecting an additional view every 100 epochs until reaching a total of 20 training views (For fewer training views, see Appendix Sec. C). We apply the same active selection strategy consistently across all datasets. For consistency, each model is initialized with the same random seed and trained for 20,000 iterations. All other settings remain unchanged across experiments, except for the view selection algorithms.
@@ 4.2 Comparison on Uncertainty Estimation
We evaluate the effectiveness of our method in improving uncertainty estimation accuracy. Specifically, we examine whether incorporating residual loss (+ResGS) and self-augmented prefiltering (+SA-ResGS) enhances the alignment between depth errors and predicted uncertainties under otherwise identical conditions. To measure this, we utilize the Area Under the Sparsification Error (AUSE) metric—a standard for evaluating uncertainty calibration adopted in (fisherrf; bayesrays; cfnerf)—where lower scores indicate better alignment between uncertainty and actual error. Thus, a lower AUSE score indicates better alignment between uncertainty predictions and actual errors, reflecting superior uncertainty calibration.
Following the approach used in CF-NeRF cfnerf, we employ depth maps from the NerfingMVS nerfingmvs network, optimized using stereo depth from COLMAP at test time. Experiments are conducted on the all nine scenes from Mip-NeRF 360 under identical view selection and evaluated using all test views.
@@ 5 Conclusion
—  Supplementary Material  —SA-ResGS: Self-Augmented Residual 3D Gaussian Splatting for Next Best View Selection
@@ A.1 Coverage Estimation
We follow the baseline setup of FisherRF [fisherrf], initializing our experiments with a sparse point cloud reconstructed via COLMAP [colmap-0, colmap-1]. This SfM output defines the axis-aligned bounding volume for discretizing the scene into voxel grids. To enable physically grounded view selection, we construct Self-Augmented Points (SA-Points) via triangulation between a training view and its extrapolated view. SA-Points are used to estimate surface occupancy and to encode observed geometry into binary voxel features. Candidate views are scored based on their voxel-level dissimilarity to the encoded training views, supporting robust coverage estimation without relying on early uncertainty signals.
@@ A.1.1 Self-Augmented Points (SA-Points) Generation
Dense Correspondence Matching. We compute dense correspondences between the ground truth image of original training view and the rendering from extrapolated view using the pretrained MASt3R model [mast3r] (MASt3R_ViTLarge_BaseDecoder_512_catmlpdpt_metric), which is robust to moderate viewpoint perturbations. As discussed in the main paper, this 3D vision foundation model produces context-aware dense correspondences by capturing structured scene semantics over 16×16 local patches, even when the extrapolated view includes geometric artifacts. This robustness allows us to extract reliable matches despite distortions in the rasterized extrapolated images, enabling consistent 3D reconstruction from single-view observations. Consequently, we can leverage extrapolated viewpoints to augment sparse training views without requiring multiview supervision.
@@ A.2.1 Dataset Detail
While this datasets provide a controlled and well-curated benchmark, its uniform view distribution limits the difficulty of next-best-view (NBV) evaluation. Such settings often make NBV strategies appear less critical, since even simple heuristics (e.g., furthest-distance selection) already perform reliably under balanced coverage [Xiao:CVPR24:NeRFDirector]
@@ Appendix E Additional Ablation Analysis
To evaluate how the choice of dense correspondence model affects our pipeline, we replaced MASt3R with $\pi^{3}$ [pi3], a recent VGGT-based matcher that predicts poses, point maps, and tracks directly from image pairs. Unlike triangulation-based MASt3R, $\pi^{3}$ provides dense 3D predictions without requiring multi-view geometry, but its outputs lie in an internal coordinate frame and thus require alignment to the global COLMAP frame via ICP.

### v2

@@ 4 Experimental Results
We evaluate our approach on two benchmark datasets: NeRF-Synthetic (Mildenhall et al., 2020), and Mip-NeRF 360 (Barron et al., 2022). While both datasets comprise scenes ranging from synthetic object-scale to real-world outdoor environments with full 360-degree coverage, its uniform and curated camera trajectories provide limited challenge for active view selection, since even simple heuristics (e.g., furthest-distance selection) already perform reliably under balanced coverage (Xiao et al., 2024). To address this limitation, we carefully curate an extended benchmark dataset including seven diverse scenes from Deep Blending (Hedman et al., 2018) and Tanks and Temples (Knapitsch et al., 2017), which introduce unbalanced view distributions and varied scene scales that better reflect practical conditions. All experiments are conducted using images at their original resolutions, and further details on dataset curation are provided in Sec. A of the Appendix.
@@
We compare our method quantitatively and qualitatively against several active 3DGS baselines which operate solely on RGB images: FisherRF (Jiang et al., 2024), ACP (Kopanas and Drettakis, 2023), and random view selection. We also include 2D-based view selection methods adopted from Active View Selector framework (Wang et al., 2025b). Following this framework, we incorporated two image quality assessment (IQA) models (MUSIQ (Ke et al., 2021) and CrossScore (Wang et al., 2024b)) to evaluate perceptual quality. Both models were re-implemented according to the authors’ official instructions and publicly available code.
@@ 4.1 Active View Selection
Table 1: Quantitative results for the Active View Selection. We compare our model with (1) Rule-based models (Random, ACP (Kopanas and Drettakis, 2023)), (2) 2D-based models (MUSIQ, CrossScore (Chen et al., 2024)), and 3D-based models (FisherRF (Jiang et al., 2024)). Results are averaged over 9 scenes from the Mip-NeRF 360, 7 scenes from NeRF-synthetic dataset and 7 additional scenes from the Deep Blending and Tanks and Temples. We conduct four trials for each scene and report average scores. For statistics for all trial please refer to the Appendix E.
@@
Following the experimental protocols outlined by Jiang et al. (2024), we adopt their prescribed initial view configurations and view selection schedules. Specifically, our experiments initiate with four uniformly distributed views, subsequently selecting an additional view every 100 epochs until reaching a total of 20 training views (For fewer training views, see Appendix Sec. C). We apply the same active selection strategy consistently across all datasets. For consistency, each model is initialized with the same random seed and trained for 20,000 iterations. All other settings remain unchanged across experiments, except for the view selection algorithms.
@@ 4.2 Comparison on Uncertainty Estimation
We evaluate the effectiveness of our method in improving uncertainty estimation accuracy. Specifically, we examine whether incorporating residual loss (+ResGS) and self-augmented prefiltering (+SA-ResGS) enhances the alignment between depth errors and predicted uncertainties under otherwise identical conditions. To measure this, we utilize the Area Under the Sparsification Error (AUSE) metric—a standard for evaluating uncertainty calibration adopted in (Jiang et al., 2024; Goli et al., 2024; Shen et al., 2022)—where lower scores indicate better alignment between uncertainty and actual error. Thus, a lower AUSE score indicates better alignment between uncertainty predictions and actual errors, reflecting superior uncertainty calibration.
Following the approach used in CF-NeRF (Shen et al., 2022), we employ depth maps from the NerfingMVS (Wei et al., 2021) network, optimized using stereo depth from COLMAP at test time. Experiments are conducted on the all nine scenes from Mip-NeRF 360 under identical view selection and evaluated using all test views.
@@ A.1 Coverage Estimation
We follow the baseline setup of FisherRF [Jiang et al., 2024], initializing our experiments with a sparse point cloud reconstructed via COLMAP [Schonberger and Frahm, 2016, Schönberger et al., 2016]. This SfM output defines the axis-aligned bounding volume for discretizing the scene into voxel grids. To enable physically grounded view selection, we construct Self-Augmented Points (SA-Points) via triangulation between a training view and its extrapolated view. SA-Points are used to estimate surface occupancy and to encode observed geometry into binary voxel features. Candidate views are scored based on their voxel-level dissimilarity to the encoded training views, supporting robust coverage estimation without relying on early uncertainty signals.
@@ A.1.1 Self-Augmented Points (SA-Points) Generation
Dense Correspondence Matching. We compute dense correspondences between the ground truth image of original training view and the rendering from extrapolated view using the pretrained MASt3R model [Leroy et al., 2024] (MASt3R_ViTLarge_BaseDecoder_512_catmlpdpt_metric), which is robust to moderate viewpoint perturbations. As discussed in the main paper, this 3D vision foundation model produces context-aware dense correspondences by capturing structured scene semantics over 16×16 local patches, even when the extrapolated view includes geometric artifacts. This robustness allows us to extract reliable matches despite distortions in the rasterized extrapolated images, enabling consistent 3D reconstruction from single-view observations. Consequently, we can leverage extrapolated viewpoints to augment sparse training views without requiring multiview supervision.
@@ A.2.1 Dataset Detail
While this datasets provide a controlled and well-curated benchmark, its uniform view distribution limits the difficulty of next-best-view (NBV) evaluation. Such settings often make NBV strategies appear less critical, since even simple heuristics (e.g., furthest-distance selection) already perform reliably under balanced coverage [Xiao et al., 2024]
@@ Appendix E Additional Ablation Analysis
To evaluate how the choice of dense correspondence model affects our pipeline, we replaced MASt3R with $\pi^{3}$ [Wang et al., 2025a], a recent VGGT-based matcher that predicts poses, point maps, and tracks directly from image pairs. Unlike triangulation-based MASt3R, $\pi^{3}$ provides dense 3D predictions without requiring multi-view geometry, but its outputs lie in an internal coordinate frame and thus require alignment to the global COLMAP frame via ICP.